OpenStudy (blaque678):

What's the next number? 1,5,9,13,17_?

OpenStudy (anonymous):

1 + 4 = 5 5 + 4 = 9 9 + 4 = 13 13 + 4 = 17 17 + 4 = 21

OpenStudy (anonymous):

see it is an Arithmetic Sequence. and the common difference is +4

OpenStudy (e.cociuba):

17, 21, 25...etc

OpenStudy (e.cociuba):

1, 5, 9, 13, 17, 21, 25, 29, 33, 37, 41, 45 and so on. Every number is 4 more than the previous number.